OpenStudy (anonymous):

I have tried this several times but can't solve it. Can you guys explain it to me? Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

Uh, did you just post the entire answer sheet to the quiz?

OpenStudy (anonymous):

'cause I don't want this, I just need someone to explain HOW to do the problem.

OpenStudy (anonymous):

@Hero can you help?

OpenStudy (anonymous):

Anyone there?

hero (hero):

Divide both sides of the second equation by 2

OpenStudy (anonymous):

OK, hang on.

OpenStudy (anonymous):

Sorry I'm a little slow, but in Substitution method don't you subtract the 2x in the first equation so you get Y = 1 - 2x? which you then substitute into the second equation??

hero (hero):

@studygirl, I was getting to that. However, there is a reason I wanted to you divide both sides of the second equation by two first.

OpenStudy (anonymous):

OK give me a minute and I'll do it. :)

OpenStudy (anonymous):

you divide 4x, 2y, and -2 all by 2? Sorry math isn't my strong suit

hero (hero):

Yes, let me know what you get afterwards.

OpenStudy (anonymous):

ok 1 second. :)

OpenStudy (anonymous):

2x + 1y = -1 ??

hero (hero):

Okay which means you are left with the following system: 2x + y = 1 2x + y = -1

OpenStudy (anonymous):

OK what do I do now?

hero (hero):

Now from that what observations can you make about the system?

OpenStudy (anonymous):

well they seem identical except that the first on is 1 and the second is -1.

hero (hero):

Yes the left sides are the same but the right side isn't. That basically tells you that this system is inconsistent (has no solution)

hero (hero):

So in essence, there is really no point in trying to solve this system. Even if you tried substitution, it wouldn't work. But if you wish to satisfy your own curiosity, you can try solving by substitution just to see what happens.

OpenStudy (anonymous):

Sorry my computer is acting up a bit. :)

OpenStudy (anonymous):

so is there no solution? or infinite solutions?

hero (hero):

You can only have infinite solutions if both equations are the same or are multiples of each other.

OpenStudy (anonymous):

so no solutions?

hero (hero):

If you had something like 2x + y = 1 4x + 2y = 2 This system would have infinite solutions because 4x + 2y = 2 is a multiple of 2x + y = 1 Try to understand why this is the case.

OpenStudy (anonymous):

OK, so I'll admit I'm a little lost as to the answer. Sorry. :(

OpenStudy (radar):

It probably would be good if you went ahead in your original direction by using substitution as instructed by the problem. You were correct using the first equation you came up with y=1-2x. Substituting this in the 2nd equation you would get: 4x +2(1-2x) = -2 then continuing 4x + 2 -4x = -2 2=-2 An inconsistency which means there is no solution. this is of course what Hero came up with by looking at the system as whole.

hero (hero):

There is no solution to your system (as I had already stated above)

OpenStudy (anonymous):

OK thanks you really are a Hero!! :) thanks for your help too Radar!

OpenStudy (anonymous):

last question that would make it the second to last answer, right?

hero (hero):

@StudyGirl21, I believe you have enough information to figure out what the answer is on your own.

OpenStudy (anonymous):

OK, I'll go with the second to last one, thanks! :)

OpenStudy (radar):

Yes, you see the x term dropped out in the process of attempting a solution by the substitution method.

OpenStudy (anonymous):

Thanks you two!! :)

OpenStudy (radar):

You're welcome and good luck in your studies. Love the profile.

OpenStudy (anonymous):

thanks! Bye! :)